# Presto! Distilling Steps and Layers for Accelerating Music Generation

### v1

@@ A.4 Analyzing Failure Modes of Combined Layer and Step Distillation
A.5 Inference-time noise schedule Sensitivity analysis
@@
A.6 RTF Analysis
We define the RTF for a model $\bm{\theta}$ as: $\text{RTF}_{b}(\bm{\theta})=\frac{bT_{\bm{\theta}}}{\text{latency}_{\bm{\theta}}(b)}$, where $T_{\bm{\theta}}$ is the generation duration or how much contiguous audio the model can generate at once and $\text{latency}_{\bm{\theta}}(b)$ is the time it takes for generating a batch of $b$ samples, following (Evans et al., 2024b; Zhu et al., 2024). This is different from the fixed-duration batched RTF used in Nistal et al. (2024). We test $b=1$ as well as the maximum batch size we could attain for each model on a single A100 40GB to get a sense of maximum throughput. We show results in Table 3 and Table 4 for all components of our generative process, including latency metrics for generation (i.e. the diffusion model or distilled generator), decoding (i.e. VAE decoder from latents to audio) and the optional mono-to-stereo (M2S), as well as overall RTF/latency for mono and stereo inference. We omit the MusicGen models and the other step-distillation methods which share the same RTF as Presto-S. For the fastest model Presto-LS, the biggest latency bottleneck is the mono-to-stereo model (Zhu et al., 2024) and VAE decoder. In future work, we hope to optimize the VAE and mono-to-stereo modules for faster inference.
@@ A.6 RTF Analysis
A.7 Presto-L Design Ablation
@@
A.8 Discrete-Time Failure Modes
@@
A.9 Listening Test Results
@@
A.10 Rejection Sampling Results
We show rejection sampling results where we generate a batch during inference and then use CLAP to reject the $r$ least similar generations to the input text prompt. CLAP rejection sampling noticeably improves CLAP Score and maintains (and sometimes improves) FAD and MMD, but reduces diversity as expected.

### v2

@@ A.7 Analyzing Failure Modes of Combined Layer and Step Distillation
A.8 Inference-time noise schedule Sensitivity analysis
@@
A.9 RTF Analysis
We define the RTF for a model $\bm{\theta}$ as: $\text{RTF}_{b}(\bm{\theta})=\frac{bT_{\bm{\theta}}}{\text{latency}_{\bm{\theta}}(b)}$, where $T_{\bm{\theta}}$ is the generation duration or how much contiguous audio the model can generate at once and $\text{latency}_{\bm{\theta}}(b)$ is the time it takes for generation following (Evans et al., 2024b; Zhu et al., 2024). This is different from the fixed-duration batched RTF used in Nistal et al. (2024). We test $b=1$ as well as the maximum batch size we could attain for each model on a single A100 40GB to get a sense of maximum throughput. We show results in Table 3 and Table 4 for all components of our generative process, including latency metrics for generation (i.e. the diffusion model or distilled generator), decoding (i.e. VAE decoder from latents to audio) and the optional mono-to-stereo (M2S), as well as overall RTF/latency for mono and stereo inference. We omit the MusicGen models and the other step-distillation methods which share the same RTF as Presto-S. For the fastest model Presto-LS, the biggest latency bottleneck is the mono-to-stereo model (Zhu et al., 2024) and VAE decoder. In future work, we hope to optimize the VAE and mono-to-stereo modules for faster inference.
@@ A.9 RTF Analysis
A.10 Presto-L Design Ablation
@@
A.11 Discrete-Time Failure Modes
@@
A.12 Listening Test Results
@@
A.13 Rejection Sampling Results
We show rejection sampling results where we generate a batch during inference and then use CLAP to reject the $r$ least similar generations to the input text prompt. CLAP rejection sampling improves CLAP Score and maintains (and sometimes improves) FAD and MMD, but reduces diversity.